OpenStudy (anonymous):

^4square root16x^2y^7 is equivalent to 1. 2x^1/2 y^7/4 2.2x^8 y^28 3. 4x^1/2 y^7/4 4.4x^8 y^28

OpenStudy (anonymous):

4 sqrt(16x^2 y^7) Protip: sqrt(abc) = sqrt(a)sqrt(b)sqrt(c). Therefore, 4sqrt(16x^2y^7) = 4sqrt(16)sqrt(x^2)sqrt(y^7) We can deal with sqrt(16); it is equal to 4. 4(4) sqrt(x^2)sqrt(y^7) 16sqrt(x^2)sqrt(y^7) Next, it is a known fact that sqrt(z^2) is the same as |z|, or the absolute value of z. That means sqrt(x^2) = |x| 16|x| sqrt(y^7) Split y^7 as y^6 * y 16|x| sqrt(y^6 * y) 16|x| sqrt(y^6) sqrt(y) 16|x| sqrt([y^3]^2) sqrt(y) 16|x| |y^3| sqrt(y)

OpenStudy (anonymous):

So what's the answer. ??